Results from the run 60705
1. General Information
RUN 60705 |
Events |
3254 |
Detectors |
SPD - SSD - TPC |
SSD DDLs |
16 |
SSD LDCs |
3 |
Duration |
08:10:22 |
The overall information for this run can be found
here
.
The number of the SPD
FastOR triggered events and the total number of golden clusters per layer are seen below:
2. OCDB files
The OCDB files that were used for the reconstruction of this run were produced by the SSD DA during the pedestal run with the number
60611. The plots below show the map of the bad channels:
3. Analysis of clusters
Here are some histograms obtained with the official offline QA for the SSD clusters. In all plots we see the accumulated statistics from all the events of this run. There was a selection on the
golden clusters.
Here is a plot of the strip occupancy for this run. The bump at 0.10-0.15 is abnormal. I guess we should compare some of the noise files around the time of this run. (Module 2050 seems to be noisy, but I'm not sure whether this one gives the golden clusters)

--
PanosChristakoglou - 25 Sep 2008
Most of the effect seems to be in Module 2170:
There are two chips that show high occupancy in this module (last chip fro P-side, first from N-side). Note also that the common-mode values for those chips look funny. What could have happened is that a strip in each chip went bad, and this screws up the common mode calculation.
Would be nice if we could catch this condition in the DA. I'm looking at the noise values from the DA now....
Run 60545 (before pedestal run 60611)
The latest pedestal run taken before run 60705 was run 60611. Looking back a global partition run (60545) before that pedestal run, I get the following occupancy plot for this module.
There are three noisy strips: 724, 801 and 802. Also known as 724P, 735N and 734N.
The noise determinantion from the DA for these strips is:
I-TAlienFile::Open: Accessing image 1 of alien:///alice/data/2008/LHC08d/OCDB/ITS/Calib/NoiseSSD/Run60611_999999999_v27_s0.root in SE <ALICE::CERN::SE>
Reading the new format of the calibration file...
iMod 1670 strip 722 noise P 2.40356 N 1.9172
iMod 1670 strip 723 noise P 2.97135 N 1.98174
iMod 1670 strip 724 noise P 2.90609 N 1.8563
iMod 1670 strip 725 noise P 2.46059 N 1.95268
iMod 1670 strip 726 noise P 2.44344 N 2.07997
iMod 1670 strip 727 noise P 2.19976 N 2.25088
iMod 1670 strip 728 noise P 2.17261 N 1.6778
iMod 1670 strip 729 noise P 2.20553 N 1.94448
iMod 1670 strip 730 noise P 2.12746 N 2.04605
iMod 1670 strip 731 noise P 2.35341 N 2.11997
iMod 1670 strip 732 noise P 2.31427 N 1.86246
iMod 1670 strip 733 noise P 2.25014 N 1.68428
iMod 1670 strip 734 noise P 2.0395 N 1.87122
iMod 1670 strip 735 noise P 2.04028 N 2.20289
iMod 1670 strip 736 noise P 2.07042 N 2.07555
iMod 1670 strip 737 noise P 1.97662 N 1.98966
iMod 1670 strip 738 noise P 2.12634 N 1.71464
So nothing suspicious.
And bad channel map made by the DA has:
I-TAlienFile::Open: Accessing image 1 of alien:///alice/data/2008/LHC08d/OCDB/ITS/Calib/BadChannelsSSD/Run60611_999999999_v26_s0.root in SE <ALICE::CERN::SE>
Reading the new format of the calibration file...
Module: 2170 Strip P: 0 - 11
Module: 2170 Strip P: 1 - 11
Module: 2170 Strip P: 2 - 11
Module: 2170 Strip P: 3 - 11
Module: 2170 Strip P: 4 - 11
Module: 2170 Strip P: 5 - 11
Module: 2170 Strip N: 9 - 11
Module: 2170 Strip P: 15 - 11
Module: 2170 Strip N: 17 - 11
Module: 2170 Strip P: 20 - 11
Module: 2170 Strip N: 21 - 11
Module: 2170 Strip P: 111 - 11
Module: 2170 Strip N: 155 - 11
Module: 2170 Strip P: 157 - 11
Module: 2170 Strip P: 432 - 11
Module: 2170 Strip N: 436 - 11
Module: 2170 Strip P: 749 - 11
Module: 2170 Strip P: 765 - 11
Module: 2170 Strip P: 766 - 11
Module: 2170 Strip N: 766 - 11
Module: 2170 Strip P: 767 - 11
New pedestal run 60788
A new pedestal run was taken today and indeed a whole bunch of new bad strips appear:
Module: 2170 Strip P: 704 - 11
Module: 2170 Strip P: 705 - 11
Module: 2170 Strip P: 707 - 11
Module: 2170 Strip N: 714 - 11
Module: 2170 Strip N: 715 - 11
Module: 2170 Strip N: 716 - 11
Module: 2170 Strip P: 723 - 11
Module: 2170 Strip N: 734 - 11
Among them is strip 723P, which is one-off (so probably the same; I just made a mistake reading from the old histo) from the one I mentioned above as being suspicious already at an earlier time. 734N, the other candidate-noisy strip is also there.
here are the noise values:
iMod 1670 strip 722 noise P 2.55345 N 2.06708
iMod 1670 strip 723 noise P 97.173 N 2.11459
iMod 1670 strip 724 noise P 15.6541 N 2.03784
iMod 1670 strip 725 noise P 3.2387 N 2.18262
iMod 1670 strip 726 noise P 6.24391 N 2.14293
iMod 1670 strip 727 noise P 2.76114 N 2.16968
iMod 1670 strip 728 noise P 2.73682 N 1.98549
iMod 1670 strip 729 noise P 2.74855 N 2.01619
iMod 1670 strip 730 noise P 2.63323 N 2.07572
iMod 1670 strip 731 noise P 2.76521 N 2.52774
iMod 1670 strip 732 noise P 2.80242 N 1.99301
iMod 1670 strip 733 noise P 2.66268 N 3.80636
iMod 1670 strip 734 noise P 2.68156 N 21.9887
iMod 1670 strip 735 noise P 2.26507 N 2.54498
So it looks like we have a clean example of two strips gone bad and it seems to be caught by the DA.
I had a quick look at run 60813, taken after this pedestal run, and it looks very clean (too clean if you ask me).
--
MarcoVanLeeuwen - 25 Sep 2008